At a given temperature, 0.00332 g of Hg in the gas phase has a pressure of 0.00120 mmHg and a volume of 435 L. What is its temperature?

Not all Gas Law problems have Kelvin (K) as the unit of temperature. They can be expressed in Celsius (°C) and Fahrenheit (°F). So, convert 123°C to K.

A 0.0997 mol sample of O2, has a pressure of 0.692 atm and a temperature of 333 K. What is its volume?
1
Expert's answer
2020-03-30T10:52:52-0400

1) Hg)=m/M=0.00332/201=0.0000165mol,

p = 0.00120 mmHg = 0.16 pa,

V = 435 L = 0.435 m3,

p*V=n*R*T

T = p*V/n*R = 0.16*0.435/0.0000165*8.31 = 507.6 K = 234.45 C


2) 123C = 396.15K


3) p = 0.692 atm = 70117 pa

V = n*R*T/p = 0.0997*8.31*333/70.117 = 3.94 m3